Selected OPR plots for some troublesome runs
- Run 26538 with noticeable depletion in phi. The logbook says that we lost part of the DCH.
- Run 26725 is BAD (rejected by Trigger) because quantities in BLT and L3 track are affected by some DCH/ROM problems.
Last modified: Fri Aug 26 10:09:01 PDT 2005